# Exploiting Unlabeled Data with Multiple Expert Teachers for Open Vocabulary Aerial Object Detection and Its Orientation Adaptation

### v1

@@ IV-C1 Horizontal Box Selection Strategies
RPN Score. The RPN Score represents the confidence that a region contains a object. This strategy filters out boxes with low foreground confidence score, which is a common approach adopted by most OVD methods [91, 23]. As Shown in Fig. 5(a), the RPN Score can be directly obtained from the region proposal network:
RPN $\text{RPN Score}=\Phi_{\text{RPN}}^{\text{score}}\circ\Phi_{\text{ENC}}(T^{\prime}(\mathbf{x}))$ (7)
Regression Jittering Variance (RJV). The RJV Score reflects the stability of object location predictions across $M$ regression times. Regression jittering means we iteratively put the predicted box into the regression branch ($\Phi_{\text{LOC}}$) for a more precise prediction. The RJV Score is defined as
$\bar{\sigma_{i}}=\frac{1}{4}\sum_{k=1}^{4}\frac{\sigma_{ik}^{2}}{(h_{-1}^{2}+w_{-1}^{2})}$ (8)
where $\{\sigma_{ik}\}_{k=1}^{4}$ is the standard derivation of the coordinates (i.e., $x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$) for the $i$-th set of regression boxes. The denominators $h_{-1}$ and $w_{-1}$ are the height and width of the last regression box, respectively, which serve to normalize the variance relative to the size of the box, thus making the score scale-invariant.
Box Jittering Variance (BJV). The BJV Score is employed to measure the stability of object location predictions across multiple jittered boxes. Box jittering involves introducing random noise around a predicted bounding box $b_{i}$ to generate $M$ jittered boxes. These jittered box are then processed by the box regression branch ($\Phi_{\text{LOC}}$) to predict the refined bounding boxes, $\{\hat{b_{i,j}}\}$, where $j$ ranges from 1 to $M$. The BJV is defined as follows
$\bar{\sigma_{i}}=\frac{1}{4}\sum_{k=1}^{4}\frac{\sigma_{ik}}{0.5(h_{i}+w_{i})}$ (9)
Here, $\{\sigma_{ik}\}_{k=1}^{4}$ represents the standard derivation of the coordinates (i.e., $x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$) for the $i$-th set of refined boxes. The terms $h_{i}$, $w_{i}$ denote the mean height and width of the $i$-th boxes set, respectively [32].
IV-C2 Hybrid Training
We follow the student-teacher training mechanism. In each training iteration, three types of data flow are sampled from the training batch: The labeled data is directly used to supervise the student ($\mathcal{L}_{s}$). The unlabeled data is assigned pseudo-labels by the localization teacher and contributes to the unsupervised loss ($\mathcal{L}_{u}$). The pseudo-labeled data from the dynamic label queue encourages the model to discover novel objects ($\mathcal{L}_{d}$), as depicted in Fig. 3. Thus, the overall loss comprises three components:
@@ IV-C2 Hybrid Training
where $\mathcal{L}_{s}$, $\mathcal{L}_{u}$ and $\mathcal{L}_{d}$ represent the supervised loss for labeled images, unsupervised loss for unlabeled images with pseudos generated by the localization teacher, and unsupervised loss for images sampled from the dynamic label queue, respectively. The weights $\alpha$, $\beta$ and $\gamma$ balance these loss components.
@@
$\mathcal{L}_{s}=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathcal{L}_{\mathrm{cls}}(\hat{s}_{i},c_{i})+\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{reg}}(\hat{b}_{i},b_{i})$ (11)
@@
$\mathcal{L}_{u}^{\mathrm{cls}}=\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{i},\hat{c}_{i}\right)+\sum_{j=1}^{N_{b}^{\mathrm{bg}}}w_{j}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{j},\hat{c}_{j}\right)$ (12)
@@
For training the box regression branch, we first apply the BJV score to filter candidates, as described in Sec. IV-B2. The regression loss is then defined as:
$\mathcal{L}_{u}^{\mathrm{reg}}=\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{reg}}\left(\hat{b}_{i}^{\mathrm{fg}},\hat{b}_{i}\right)$ (13)
@@
Queue Data Flow. To enhance the open-vocabulary capability, we randomly sample a batch of images from the dynamic label queue to train the student, see Sec. IV-B3. Like unsupervised loss, $\mathcal{L}_{d}$ also includes two parts: $\mathcal{L}_{d}^{\mathrm{cls}}$ and $\mathcal{L}_{d}^{\mathrm{reg}}$. Since the localization teacher primarily guides the student in discovering and localizing objects, the $\mathcal{L}_{d}^{\mathrm{reg}}$ is optional in this process. The classification loss is defined as follows:
$\mathcal{L}_{d}^{\mathrm{cls}}=\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{i},\hat{c}_{i}\right)+\sum_{j=1}^{N_{b}^{\mathrm{bg}}}w_{j}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{j},\hat{c}_{j}\right)$ (14)
@@
IV-D Oriented Open-vocabulary Detection
Aerial objects often exhibit arbitrary orientation, the generic horizontal object detector cannot locate these oriented aerial objects accurately. Therefore, we develop oriented open-vocabulary detector to address this problem. We utilize the Oriented R-CNN [27] framework as the base detector, which is capable of representing the objects with their oriented bounding boxes, each box is denoted by a 5-dimensional vector $b_{i}=(c_{x},c_{y},w,h,a)$, where $(c_{x},c_{y})$ is the center coordinate of the box, $w$ and $h$ represent the width and height of the box, and $a$ specifies the angle of rotation. The training objective are almost the same as horizontal detection training, as detailed in Sec. IV-C2. However, when generating oriented pseudo-labels, it is crucial to consider both the scale and orientation of the bounding boxes. This section describes two advanced strategies for selecting robust pseudo oriented boxes.
IV-D1 Oriented Box Selection Strategies
@@
Box Scale Jittering Variance (SJV). The SJV Score evaluates the scale stability across multiple noise instances. We first randomly add noise $(\Delta w,\Delta h,\Delta a)$ to box representations $(w,h,a)$, thus generate $M$ jittered boxes. Here, we fix the center point $(c_{x},c_{y})$ for each box. Then, the noised boxes are refined by the oriented regression branch. Finally, the scale jittering variance is calculated by
$\bar{\sigma_{i}}=\frac{1}{2}\sum_{k=1}^{2}\frac{\sigma_{ik}}{0.5(h_{i}+w_{i})}$ (15)
Here, $\sigma_{ik}$ represents the standard deviation of the scale components (height and width), $h_{i}$ and $w_{i}$ are the mean height and width of the $i$-th set of jittered boxes.
@@
$\bar{\sigma_{i}}=\mathrm{Var}(\mathrm{f}(a_{i1}),\mathrm{f}(a_{i2}),\cdots,\mathrm{f}(a_{ik}))$ (16)
where $a_{ij}$ denotes the angle of $j$-th box of $i$-th box set. The function $\mathbf{f}(\cdot)$ denotes a function of the angle, which transform the angle values to better capture the variance in orientation, e.g., $\sin(\cdot)$ or $\text{Identity}(\cdot)$ function.
V Experiments
V-A Open-vocabulary Aerial Detection Benchmark
As shown in Fig. 6, we evaluate our approach and set the open-vocabulary aerial detection benchmark [^2] The code is available at https://github.com/lizzy8587/CastDet. on two typical aerial datasets: DIOR and DOTA. Following the setup of CastDet [36], we divide the classes into base ($\mathcal{C}_{\text{base}}$) and novel ($\mathcal{C}_{\text{novel}}$) categories. Specifically, we use 16 base classes and 4 novel classes for DIOR, 13 base classes and 2 novel classes for DOTA. During training, only the annotations for base categories are used. Our experiments cover two tasks: horizontal open-vocabulary detection and oriented open-vocabulary detection.
V-A1 Datasets.
@@ V-A Open-vocabulary Aerial Detection Benchmark
DOTA [12] comprises 2,806 aerial images from different sensors and platforms with the resolution range from 800 to 4,000, annotated by 15 common categories. We crop the original images into 1,024$\times$1,024 patches with an overlap of 200, and follow the guidelines in MMRotate [92] to merge the detection results of all patches for evaluation.
@@
DIOR [11] provides a large collection of 23,463 images, each with a resolution of $800\times 800$. These images encompass 192,472 instances across 20 object classes. We extract 8,725 images to serve as unlabeled data. Evaluation is conducted on the test set.
@@
VisDroneZSD [25] is a subset of DIOR [11], which contains 8,730 images for training and 3,337 images for testing. We follow the generalized zero-shot detection (GZSD) setting in VisDrone2023 Challenge to split the base/novel categories and train/test datasets of DIOR.
@@
NWPU-RESISC45 [93] is a large-scale remote sensing classification dataset, covering 45 scene classes with 700 images in each class. NWPU-RESISC45 is used as extra weakly supervised training data for comparison experiment.
V-A2 Algorithms.
@@
For the open-vocabulary aerial detection task, we establish two types of OVAD benchmarks: HBB-based and OBB-based open-vocabulary detectors. Due to the absence of oriented open-vocabulary detection algorithms, we integrate orientation prediction into several well-established open-vocabulary detection frameworks. We involve significant improvements to these popular frameworks: ViLD [34], GLIP [19], and GroundingDINO [21]. These models are originally designed for horizontal object detection, which are adeptly transformed to predict the orientation of objects, resulting in the development of Oriented ViLD, Oriented GLIP, and Oriented GroundingDINO.
Figure 7: Visualization of RPN foreground confidence during the training process. S, LT, ET denote student, localization teacher and external teacher, respectively.
@@
TABLE I: Effectiveness of hybrid training in enhancing novel category discovery. A comparative analysis of region proposal recall for different training strategies. S, LT, ET denote the student, localization teacher and external teacher, respectively.
@@
V-A3 Evaluation Metrics.
In our evaluation, we adopt mean Average Precision (mAP), mean Average Recall (mAR), and Harmonic Mean (HM) as key metrics. The mAP and mAR metrics are computed at an Intersection over Union (IoU) threshold of 0.5. To provide a comprehensive evaluation, we follow the approach proposed in [25] and incorporate the Harmonic Mean. This metric provides a balanced perspective on overall performance, by evaluating results across base and novel categories. It can be applied to both mAP and mAR metrics, for example, using mAP as an illustration, HM is defined as follows:
$\mathrm{HM}=2\frac{{\mathrm{mAP_{base}}}\cdot{\mathrm{mAP_{novel}}}}{{\mathrm{mAP_{base}}}+{\mathrm{mAP_{novel}}}}$ (17)
V-A4 Implementation Details.
@@
V-B Ablation studies
Hybrid Training
To demonstrate the effectiveness of hybrid training in facilitating the discovery of novel categories, we evaluate the RPN recall of the detector across three training paradigms: supervised training (S), closed-vocabulary student-teacher semi-supervised training (S+LT), and our open-vocabulary hybrid training (S+LT+ET). For fairness, we employ class-agnostic recall, i.e., any object proposed by the RPN is considered successfully detected. As presented in Table I, our hybrid training strategy yields substantial improvements in RPN recall for novel categories compared to the semi-supervised approach (e.g., 69.1 vs. 47.7 mARnovel).
H-Box Selection Strategy
We evaluate the effectiveness of different horizontal box selection metrics, including the RPN Score, RJV, and BJV, with the results summarized in Table II. The regression jittering strategy achieves a superior performance of 47.8% mAPnovel. As shown in Fig. 9(a)$\sim$(c), regression jittering exhibits a stronger correlation with both the classification score and the IoU score, enabling the selection of more accurate pseudo-labels, which enhances the training process. The qualitative results of pseudo-labels under different box selection strategies are presented in Fig. 8.
TABLE II: Comparison on horizontal box selection strategies.
@@ V-B Ablation studies
TABLE III: Comparison on oriented box selection strategies.
Dataset Strategy mAP mAPbase mAPnovel HM
DIOR RPN Score 45.7 51.1 23.9 32.6
AJV Score 45.7 51.1 24.2 32.8
SJV Score 45.9 51.4 23.9 32.6
AJV+SJV 45.9 51.3 24.3 33.0
DOTA RPN Score 57.6 62.2 27.6 38.3
AJV Score 57.4 61.6 30.2 40.6
SJV Score 57.2 61.1 31.6 41.7
AJV+SJV 58.1 61.7 34.9 44.6
TABLE IV: Ablation on dynamic label queue.
Dynamic Queue mAP mAPbase mAPnovel HM
11.6 9.7 19.3 12.9
37.7 36.3 43.5 39.6
38.1 37.2 41.5 39.2
39.5 38.6 43.3 40.8
@@
We compare the effects of two oriented box selection strategies: SJV and AJV. Fig.9(d)$\sim$(e) illustrates improved score correlation when employing SJV and AJV, which enhances the model training. Furthermore, Table III shows that using both AJV and SJV score yields better results, e.g., achieving 24.3% mAPnovel on DIOR-R and 34.9% mAPnovel on DOTA-R. The visualization of selected pseudo-boxes is shown in Fig. 8.
@@
To demonstrate the effectiveness of dynamic label queue, we conduct experiments on whether to adopt a dynamic update strategy or whether to use a label queue for storing pseudo-labels Table IV shows substantial improvement achieved through applying dynamic updating strategy or using a label queue.
@@
Figure 10: Precision and recall statistics for novel categories of pseudo-labels in dynamic label queue. (a) Precision-recall curve for different iterations. (b) The mAP and mAR curve on iterations. (c) Visualization of pseudo-labels across different iterations.
@@
To demonstrate that our model is well suited for the aerial scenario where labeled data is limited, we conduct label fraction experiments with 34%, 50% and full labeled data. Table V shows that even with only 34% of the original labeled data, the performance remains comparable, with only a minor reduction of 0.9% mAP.
TABLE V: Ablation on fraction of labeled data.
@@
Table VI shows that $\alpha=1.0$, $\beta=2.0$ and $\gamma=1.0$ could be better for novel categories.
TABLE VI: Ablation with different loss weights defined in Equ. (10).
@@
TABLE VII: Ablation with Different Background Embeddings.

### v2

@@ 3.3.1 Horizontal Box Selection Strategies
RPN Score. The RPN Score represents the confidence that a region contains a object. This strategy filters out boxes with low foreground confidence score, which is a common approach adopted by most OVD methods [zhao2022exploiting, chen2023ovarnet]. As Shown in Fig. 7(a), the RPN Score can be directly obtained from the region proposal network:
$\text{RPN Score}=\Phi_{\text{RPN}}^{\text{score}}\circ\Phi_{\text{ENC}}(T^{\prime}(\mathbf{x})).$ (7)
Regression Jittering Variance (RJV). The RJV Score reflects the stability of object location predictions across $M$ regression steps. To achieve a more precise localization, we employ regression jittering, which iteratively refines the predicted box by feeding it back into the regression branch ($\Phi_{\text{LOC}}$) for multiple steps. Specifically, given an initial predicted box, we iteratively pass this box through the regression branch $M$ times. Each iteration outputs a refined bounding box prediction, where slight variations in the coordinates may occur. The RJV Score quantifies this instability by measuring the variance of the predicted coordinates across these regression steps. Formally, the RJV Score is defined as
$\bar{\sigma_{i}}=\frac{1}{4}\sum_{k=1}^{4}\frac{\sigma_{ik}^{2}}{(h_{-1}^{2}+w_{-1}^{2})},$ (8)
where $\{\sigma_{ik}\}_{k=1}^{4}$ represents the standard derivation of the four box coordinates (i.e., $x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$) for the $i$-th set of regression boxes obtained across $M$ iterations. The denominators $h_{-1}$ and $w_{-1}$ denote the height and width of the final predicted box after $M$ steps, which serve to normalize the variance relative to the size of the box, thus making the score scale-invariant.
Box Jittering Variance (BJV). The BJV Score is employed to measure the stability of object location predictions across multiple jittered boxes. Box jittering involves introducing random noise around a predicted bounding box $b_{i}$ to generate $M$ jittered boxes. These jittered box are then processed by the box regression branch ($\Phi_{\text{LOC}}$) to predict the refined bounding boxes $\{\hat{b_{i,j}}\}$, where $j$ ranges from 1 to $M$. The BJV is defined as follows
$\bar{\sigma_{i}}=\frac{1}{4}\sum_{k=1}^{4}\frac{\sigma_{ik}}{0.5(h_{i}+w_{i})},$ (9)
where $\{\sigma_{ik}\}_{k=1}^{4}$ represents the standard derivation of the coordinates (i.e., $x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$) for the $i$-th set of refined boxes. The terms $h_{i}$, $w_{i}$ denote the mean height and width of the $i$-th boxes set, respectively [softteacher_xu2021end].
3.3.2 Hybrid Training
We follow the student-teacher training mechanism. In each training iteration, three types of data flow are sampled from the training batch: The labeled data is directly used to supervise the student model ($\mathcal{L}_{s}$). The unlabeled data is assigned pseudo-labels by the localization teacher model and contributes to the unsupervised loss ($\mathcal{L}_{u}$). The pseudo-labeled data from the dynamic label queue encourages the model to discover novel objects ($\mathcal{L}_{d}$), as depicted in Fig. 4. Thus, the overall loss comprises three components:
@@ 3.3.2 Hybrid Training
where $\mathcal{L}_{s}$, $\mathcal{L}_{u}$ and $\mathcal{L}_{d}$ represent the supervised loss for labeled images, unsupervised loss for unlabeled images with pseudos generated by the localization teacher model, and unsupervised loss for images sampled from the dynamic label queue, respectively. The weights $\alpha$, $\beta$ and $\gamma$ balance these loss components.
@@
$\mathcal{L}_{s}=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathcal{L}_{\mathrm{cls}}(\hat{s}_{i},c_{i})+\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{reg}}(\hat{b}_{i},b_{i}),$ (11)
@@
$\mathcal{L}_{u}^{\mathrm{cls}}=\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{i},\hat{c}_{i}\right)+\sum_{j=1}^{N_{b}^{\mathrm{bg}}}w_{j}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{j},\hat{c}_{j}\right),$ (12)
@@
For training the box regression branch, we first apply the BJV score to filter candidates, as described in Sec. 3.2.2. The regression loss is then defined as:
$\mathcal{L}_{u}^{\mathrm{reg}}=\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{reg}}\left(\hat{b}_{i}^{\mathrm{fg}},\hat{b}_{i}\right),$ (13)
@@
Queue Data Flow. To enhance the open-vocabulary capability, we randomly sample a batch of images from the dynamic label queue to train the student model, see Sec. 3.2.3. Like unsupervised loss, $\mathcal{L}_{d}$ also includes two parts: $\mathcal{L}_{d}^{\mathrm{cls}}$ and $\mathcal{L}_{d}^{\mathrm{reg}}$. Since the localization teacher model primarily guides the student model in discovering and localizing objects, the $\mathcal{L}_{d}^{\mathrm{reg}}$ is optional in this process. The classification loss is defined as follows:
$\mathcal{L}_{d}^{\mathrm{cls}}=\frac{1}{N_{b}^{\mathrm{fg}}}\sum_{i=1}^{N_{b}^{\mathrm{fg}}}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{i},\hat{c}_{i}\right)+\sum_{j=1}^{N_{b}^{\mathrm{bg}}}w_{j}\mathcal{L}_{\mathrm{cls}}\left(\hat{s}_{j},\hat{c}_{j}\right),$ (14)
@@
3.4 Oriented Open-vocabulary Detection
Aerial objects often exhibit arbitrary orientation, the generic horizontal object detector cannot locate these oriented aerial objects accurately. Therefore, we develop oriented open-vocabulary detector to address this problem. We utilize the Oriented R-CNN [orientedrcnn_Xie_2021_ICCV] framework as the base detector, which is capable of representing the objects with their oriented bounding boxes, each box is denoted by a 5-dimensional vector $b_{i}=(c_{x},c_{y},w,h,a)$, where $(c_{x},c_{y})$ is the center coordinate of the box, $w$ and $h$ represent the width and height of the box, and $a$ specifies the angle of rotation. The training objective are almost the same as horizontal detection training, as detailed in Sec. 3.3.2. However, when generating oriented pseudo-labels, it is crucial to consider both the scale and orientation of the bounding boxes. This section describes two advanced strategies for selecting robust pseudo oriented boxes.
3.4.1 Oriented Box Selection Strategies
@@
Box Scale Jittering Variance (SJV). The SJV Score evaluates the scale stability across multiple noised instances. We first randomly add noise $(\Delta w,\Delta h,\Delta a)$ to box representations $(w,h,a)$, thus generate $M$ jittered boxes. Here, we fix the center point $(c_{x},c_{y})$ for each box. Then, the noised boxes are refined by the oriented regression branch. Finally, the scale jittering variance is calculated by
$\bar{\sigma_{i}}=\frac{1}{2}\sum_{k=1}^{2}\frac{\sigma_{ik}}{0.5(h_{i}+w_{i})},$ (15)
where $\sigma_{ik}$ represents the standard deviation of the scale components (height and width), $h_{i}$ and $w_{i}$ are the mean height and width of the $i$-th set of jittered boxes.
@@
$\bar{\sigma_{i}}=\mathrm{Var}(\mathrm{f}(a_{i1}),\mathrm{f}(a_{i2}),\cdots,\mathrm{f}(a_{ik})),$ (16)
where $a_{ij}$ denotes the angle of $j$-th box of $i$-th box set. The function $\mathbf{f}(\cdot)$ denotes a function of the angle, which transforms the angle values to better capture the variance in orientation, e.g., $\sin(\cdot)$ or $\text{Identity}(\cdot)$ function. Here, $\mathrm{Var}(\cdot)$ denotes the variance of these input values.
4 Experiments
4.1 Open-vocabulary Aerial Detection Benchmark
As shown in Fig. 8, we evaluate our approach and set the open-vocabulary aerial detection benchmark [^2] The code is available at https://github.com/VisionXLab/CastDet. on three typical aerial datasets: DIOR [dior_li2020object], DOTA [DOTA8578516] and STAR [li2024star]. Following the setup of CastDet [castdet], we divide the classes into base ($\mathcal{C}_{\text{base}}$) and novel ($\mathcal{C}_{\text{novel}}$) categories. Specifically, we use 16 base classes and 4 novel classes for DIOR, 13 base classes and 2 novel classes for DOTA, and 40 base classes and 8 novel classes for STAR. During training, only the annotations for base categories are used. Our experiments cover two tasks: horizontal open-vocabulary detection and oriented open-vocabulary detection.
4.1.1 Datasets
@@
DOTA [DOTA8578516] comprises 2,806 aerial images from different sensors and platforms with the resolution range from 800 to 4,000, annotated by 15 common categories. We crop the original images into $\times1,024$ patches with an overlap of 200, and follow the guidelines in MMRotate [zhou2022mmrotate] to merge the detection results of all patches for evaluation.
@@
DIOR [dior_li2020object] provides a large collection of 23,463 images, each with a resolution of $800\times 800$. These images encompass 192,472 instances across 20 object classes. We extract 8,725 images to serve as unlabeled data. Evaluation is conducted on the test set.
STAR [li2024star] is a large-scale satellite dataset comprising very-high-resolution (VHR) images with sizes ranging from $512\times 768$ to $27,860\times 31,096$ pixels. It includes 48 object categories and over 210k instances. Following the preprocessing pipeline in MMRotate [zhou2022mmrotate], the images are cropped into $1,024\times 1,024$ patches with an overlap of 200 for training and evaluation.
@@
VisDroneZSD [VisDrone2023] is a subset of DIOR [dior_li2020object], which contains 8,730 images for training and 3,337 images for testing. We follow the generalized zero-shot detection (GZSD) setting in VisDrone2023 Challenge to split the base/novel categories and train/test datasets of DIOR.
@@
NWPU-RESISC45 [NWPU_RESISC45_7891544] is a large-scale remote sensing classification dataset, covering 45 scene classes with 700 images in each class. NWPU-RESISC45 is used as extra weakly supervised training data for comparison experiment.
4.1.2 Algorithms
@@
For the open-vocabulary aerial detection task, we establish two types of OVAD benchmarks: HBB-based and OBB-based open-vocabulary detectors. Due to the absence of oriented open-vocabulary detection algorithms, we integrate orientation prediction into several well-established open-vocabulary detection frameworks. We involve significant improvements to these popular frameworks: ViLD [gu2021open], GLIP [9879567glip], and GroundingDINO [liu2023groundingdino]. These models are originally designed for horizontal object detection, which are adeptly transformed to predict the orientation of objects, resulting in the development of Oriented ViLD, Oriented GLIP, and Oriented GroundingDINO.
4.1.3 Evaluation Metrics
In our evaluation, we adopt mean Average Precision (mAP), mean Average Recall (mAR), and Harmonic Mean (HM) as key metrics. The mAP and mAR metrics are computed at an Intersection over Union (IoU) threshold of 0.5. To provide a comprehensive evaluation, we follow the approach proposed in [VisDrone2023] and incorporate the Harmonic Mean. This metric provides a balanced perspective on overall performance, by evaluating results across base and novel categories. It can be applied to both mAP and mAR metrics, for example, using mAP as an illustration, HM is defined as follows:
$\mathrm{HM}=2\frac{{\mathrm{mAP_{base}}}\cdot{\mathrm{mAP_{novel}}}}{{\mathrm{mAP_{base}}}+{\mathrm{mAP_{novel}}}}.$ (17)
4.1.4 Implementation Details
We implement our method with MMDetection 3.3.0 [mmdetection] and MMRotate 1.0.0 [zhou2022mmrotate] toolbox. For horizontal open-vocabulary detection, we employ Faster R-CNN [ren2015faster] with ResNet50-C4 [he2016deep] backbone as the base detector. The model is initialized by a pre-trained Soft Teacher [softteacher_xu2021end] and an external teacher (e.g., RemoteCLIP [liu2023remoteclip], SkyCLIP [wang2024skyscript] or GeoRSCLIP [zhang2024georsclip]), followed by 10kk iterations of training with a batch size of 12. For oriented object open-vocabulary detection, we utilize Oriented R-CNN [orientedrcnn_Xie_2021_ICCV] as the base detector. The model is initialize with a pre-trained Oriented R-CNN [orientedrcnn_Xie_2021_ICCV] and the external teacher, followed by 10k iterations of training with a batch size of 8. We adopt Stochastic Gradient Descent (SGD) as the optimizer, using a learning rate of 0.01, with momentum and weight decay parameters set to 0.9 and 0.0001, respectively. To maintain high-quanlity pseudo-labels, we set the RPN foreground threshold and prediction probability threshold $p_{0}$ to 0.95 and 0.8, respectively, based on the analysis in Fig. 10, to filter out poor predictions.
4.2 Ablation studies
Hybrid Training
To demonstrate the effectiveness of hybrid training in facilitating the discovery of novel categories, we evaluate the RPN recall of the detector across three training paradigms: supervised training (S), closed-vocabulary student-teacher semi-supervised training (S+LT), and our open-vocabulary hybrid training (S+LT+ET). For fairness, we employ class-agnostic recall, i.e., any object proposed by the RPN is considered successfully detected. As presented in Table 1, our hybrid training strategy yields substantial improvements in RPN recall for novel categories compared to the semi-supervised approach (e.g., 69.1 vs. 47.7 mARnovel).
@@ 4.2 Ablation studies
Table 1: Effectiveness of hybrid training in enhancing novel category discovery on VisDroneZSD. A comparative analysis of region proposal recall for different training strategies. S, LT and ET represent the student model, localization teacher model and external teacher model, respectively.
@@
Figure 9: Visualization of RPN foreground confidence during the training process. S, LT and ET denote the student model, localization teacher model and external teacher model, respectively.
H-Box Selection Strategy
@@
Table 2: Comparison on horizontal box selection strategies on VisDroneZSD.
@@
[…]
Table 3: Comparison on oriented box selection strategies.
Strategy DIOR DOTA
mAP mAPbase mAPnovel HM mAP mAPbase mAPnovel HM
RPN Score 45.7 51.1 23.9 32.6 57.6 62.2 27.6 38.3
AJV Score 45.7 51.1 24.2 32.8 57.4 61.6 30.2 40.6
SJV Score 45.9 51.4 23.9 32.6 57.2 61.1 31.6 41.7
AJV+SJV 45.9 51.3 24.3 33.0 58.1 61.7 34.9 44.6
@@
To demonstrate the effectiveness of dynamic label queue, we conduct experiments on whether to adopt a dynamic update strategy or whether to use a label queue for storing pseudo-labels. Specifically, we consider the following settings: i) training the model without a dynamic queue; ii) incorporating teacher models to generate pseudo-labels at each iteration, and directly using these labels to train the student model without storing them in a queue; iii) constructing a static pseudo queue; and iv) implementing the proposed dynamic label queue. As shown in Table 4, adopting either a dynamic update strategy or a label queue leads to substantial performance improvements, with the proposed dynamic label queue achieving the most significant gains.
@@
Figure 12: Statistics and visualizations for novel categories of pseudo-labels in dynamic label queue. (a) Precision-recall curve for different iterations. (b) The mAP and mAR curve on iterations. (c) Visualization of pseudo-labels across different iterations.
Table 4: Ablation on dynamic label queue on VisDroneZSD.
ID Dynamic Queue mAP mAPbase mAPnovel HM
i 11.6 9.7 19.3 12.9
$i$ 37.7 36.3 43.5 39.6
iii 38.1 37.2 41.5 39.2
iv 39.5 38.6 43.3 40.8
Cross Evaluation
Some popular detectors for natural images, such as GroundingDINO [liu2023groundingdino] and GLIP [9879567glip], underperform on novel aerial categories due to a lack of a robust strategy for discovering novel aerial objects. Thus, we provide them with pseudo-labels filtered by our box selection strategies, i.e., AJV and RJV. As shown in Table 5, this leads to significant improvements, e.g., 6.6% mAPnovel gains for GLIP and 11.8% mAPnovel gains for GroundingDINO, highlighting the effectiveness of our approach in bridging the gap between general object detection frameworks and the specialized requirements of aerial imagery.
Table 5: Boost GLIP/GroundingDINO with our pseudo-labeling strategies on DOTA. $\ddagger$: Use pseudo-labels generated by our pseudo-labeling strategies to train the model.
Method mAP mAPbase mAPnovel HM
GLIP 57.5 65.2 8.0 14.2
GroundingDINO 58.0 66.0 5.8 10.6
GLIP 59.7 66.6 14.6 24.0
GroundingDINO 47.0 51.5 17.6 26.3
External Teacher
We evaluate different external teachers by varying both their backbone architectures (e.g., R50 vs. ViT-L-14) and vision–language models (e.g., RemoteCLIP [liu2023remoteclip], GeoRSCLIP [zhang2024georsclip], and SkyCLIP [wang2024skyscript]). As shown in Table 6, CastDet maintains strong overall performance across different choices of external teachers, with stronger architectures (e.g., SkyCLIP and ViT-L-14 backbone) achieving better performance on novel categories.
Table 6: Ablation on external teachers.
Dataset External Teacher Backbone mAP mAPbase mAPnovel HM
VisDroneZSD RemoteCLIP [liu2023remoteclip] R50 40.5 39.0 46.3 42.3
RemoteCLIP [liu2023remoteclip] VIT-L-14 40.8 38.2 51.2 43.8
GeoRSCLIP [zhang2024georsclip] VIT-B-32 41.8 40.8 45.6 43.1
GeoRSCLIP [zhang2024georsclip] VIT-L-14 41.1 39.4 48.2 43.3
SkyCLIP [wang2024skyscript] VIT-L-14 41.2 38.8 51.0 44.0
DIOR RemoteCLIP [liu2023remoteclip] R50 61.6 65.7 45.2 53.6
RemoteCLIP [liu2023remoteclip] VIT-L-14 61.6 64.4 50.5 56.6
GeoRSCLIP [zhang2024georsclip] VIT-B-32 62.4 66.6 45.2 53.9
GeoRSCLIP [zhang2024georsclip] VIT-L-14 61.6 65.1 47.6 55.0
SkyCLIP [wang2024skyscript] VIT-L-14 61.6 64.5 50.1 56.4
@@
To demonstrate that our model is well suited for the aerial scenario where labeled data is limited, we conduct label fraction experiments with 34%, 50% and full labeled data. Table 7 shows that even with only 34% of the original labeled data, the performance remains comparable, with only a minor reduction of 0.9% mAP.
Table 7: Ablation on fraction of labeled data on VisDroneZSD.
@@
Table 8 presents the performance comparison with different loss weight configurations. We observe that setting $\alpha=1.0$, $\beta=2.0$ and $\gamma=1.0$ could be better for novel categories.
Table 8: Ablation with different loss weights defined in Equ. (10) on DOTA.
@@
Background Embeddings
We compare three common background embeddings: fixed all-zero embedding [zhong2022regionclip], fixed mean embedding [fixMean_rahman2018zero], and learnable background embedding [learnBG_zheng2020background]. As shown in Table 9, the choice of background embedding has slight impact on our experimental results, with the learnable embedding showing better performance.
Table 9: Ablation with Different Background Embeddings on DOTA.